(1/4 + 1/8) / (1/2 + 1/5)
does this equal:
15/28

Answer by DrBeeee(684)

Yes
use cross multiply
1/4 + 1/8 = 12/32
1/2 + 1/5 = 7/10
(12/32)/(7/10) = 12/32 * 10/7 = 3/8*10/7 =30/56 = 15/28